# Measuring eccentricity and gas-induced perturbation from gravitational waves of LISA massive black hole binaries

### v1

@@ 9 Conclusion
In this paper, we considered GWs from eccentric MBHBs embedded in a CBD to estimate if both eccentricity and gas imprints (migration and accretion) could be concurrently measured from the emitted gravitational waves observed by LISA. We study systems of interest at redshift $z=1$ with highly spinning ($\chi_{1,2}=0.9$) individual BH masses, $M_{z}$, between $10^{4}$$10^{6}~{}{\rm M}_{\sun}$ and a primary-to-secondary mass ratio, $q\in[1.2,8]$, such that the MBHBs spend at least four years in the LISA band before merging. We considered both one-year and four-year times of coalescence ($t_{c}$) to study the measurability of the parameters for systems with initial eccentricity, $e_{0}$, from $10^{-3}$$10^{-1}$ and effective gas amplitude, $C_{\rm g}$, between $-10^{3}$ and $10^{3}$. We assumed that the power-law ($n_{\rm g}$) scaling of the environmental effect with semi-major axis is $n_{\rm g}=4$, as expected from both migration and accretion (see Sec. 3). Due to the high expected SNRs, 150similar-toabsent150$\sim$ 150∼ 1502500250025002500, we found that LISA observations should be able to place constraints on the intrinsic-inspiral variables, $e_{0}$ and $C_{\rm g}$, as well as on the intrinsic-merger binary parameters, $\{M_{z},q,\chi_{1,2},t_{c}\}$. To account for LISA’s motion around the Sun and model the time delay interferometry response, we used the lisabeta software, and included the dephasing due to gas described by Eq. (11) into the TaylorF2Ecc waveform model. We surveyed the parameter space analytically using the Fisher formalism and then studied a few cases using Bayesian inference. Finally, we assessed whether a weak environmental imprint could be confused with a small eccentric signal using GW data alone. We itemize our main findings below.
@@
  • The cross terms between eccentricity and gas are negligible (see Appendix C). Therefore, when constraining $C_{\rm g}$ and $e_{0}$ simultaneously, the relative uncertainties on $e_{0}$ and $C_{\rm g}$ are independent of their exact magnitude (see Fig. 2).
@@
  • $C_{\rm g}\gtrsim 100$ is constrained to $<50$ per cent relative error for circular binaries and $C_{\rm g}\gtrsim 1000$ for eccentric systems. This translates to confidently measuring ${\rm f}_{\rm Edd}$ to around $0.1$ and $1$ respectively in a four-year observation window for a $M_{z}=10^{6}~{}{\rm M}_{\sun}$ and $q=8.0$ MBHB embedded in an extremely thin disc (see Fig. 3 and the connection between $C_{\rm g}$ and accretion properties in Eq. 12).
@@
  • Bayesian inference verifies the results of the Fisher results, showing posteriors that overlapped with the prediction from the Fisher matrix (see Fig. 5 and that peaked at the same value of $C_{\rm g}$ with or without eccentricity (see Fig. 6).
@@
  • An eccentric vacuum template can mimic a circular environmental signal for a weak injected gas amplitude, $|C_{g,\rm inj}|\lesssim 10$. However, vice versa does not hold for small $e_{0,\rm inj}\gtrsim 10^{-2.5}$ (see Tables 3 and 4).
  • An injected eccentric environmental signal could be confused with a vacuum eccentric signal for $|C_{g,\rm inj}|\lesssim 100$ and $e_{0,\rm inj}=0.01$. Similarly, the same signal could be mimicked by a circular gas-perturbed template even for $|C_{g,\rm inj}|\sim 10^{3}$ (see Tables 5 and 6).
@@ Acknowledgements
MG, AD, and LM acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092. ST is supported by the SNSF Ambizione Grant Number: PZ00P2-202204. We acknowledge John G. Baker and Sylvain Marsat for providing us access to their lisabeta software. The authors also acknowledge use of the Mathematica software (Wolfram Research Inc., 2021), NumPy (Harris et al., 2020), and inspiration drawn from the GWFAST package (Iacovelli et al., 2022) regarding the python implementation of TaylorF2Ecc.
@@ Appendix A Definition of different terms
$\eta$
Symmetric mass ratio $q/(1+q)^{2}$
$v$
Characteristic velocity $(GM_{z}\pi f/c^{3})^{\frac{1}{3}}$
$\psi^{(0)}_{\rm TF2}$
leading-order circular phase contribution $(3/128\eta)v^{-5}$
$\Delta\psi_{\rm gas}$
leading-order gas-induced phasing correction $10^{-16}C_{\rm g}\psi^{(0)}_{\rm TF2}v^{-2n_{\rm g}}$
Intrinsic-merger parameters
$\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$
Intrinsic-inspiral parameters
$\{e_{0},C_{\rm g},n_{\rm g}\}$
Extrinsic parameters
$\{D_{\rm L},\imath,\phi_{c},\lambda,\beta,\psi\}$
Fiducial parameters in the LISA frame
$M_{z}=10^{5}~{}{\rm M}_{\sun},q=8,\chi_{1,2}=0.9,t_{c}=4~{}{\rm years},e_{0}=0.1,C_{\rm g}=100,n_{\rm g}=4,D_{\rm L}=6791.3~{}{\rm Mpc},\{\imath,\phi_{c},\lambda,\beta,\psi\}=0.5~{}{\rm radians}$
@@ Appendix B Gas-eccentricity cross terms?
Both $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are suppressed by powers of $e_{0}(v_{0}/v)^{19/6}$ and multiplied by ${\mathcal{A}}$, which is itself extremely small ($\sim\mathcal{O}(10^{-16})$) as per Eq. (19). Therefore, measurements of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ either requires them to be extremely large, or we need to observe the signal at a much lower frequency than what LISA can measure. Current simulations suggest that $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are at most $\mathcal{O}(10^{2})$ as per Table 8.151515In this work we focus only on the orbital-averaged value of the torque and neglect fluctuations that can have a higher magnitude (Zwick et al., 2022). Hence, we drop the cross-terms in Eq. (10) with $e_{0}(v_{0}/v)^{19/6}$ to recover the circular limit in Eq. (10).
Appendix C $\Delta\psi_{\rm gas}$ for a CBD model
C.1 Migration
Assuming a thin CBD torque fiducial model, we have in the circular limit:
$\bar{\Gamma}_{\rm CBD}$ $=\xi\dot{M}a^{2}\Omega=\frac{d}{dt}(\eta M_{z}a^{2}\Omega)\implies\dot{\bar{a}}_{\rm mig}=\xi\frac{\dot{M}}{M_{z}}\frac{2}{\eta}a,$
where in the first line we have ignored the mass accretion term when taking the time-derivative (see Section 3.2). Therefore, we have
$\dot{a}_{\text{mig}}$ $={\mathcal{A}}\left(\frac{a}{GM_{z}/c^{2}}\right)^{n_{\rm g}}\dot{\bar{a}}_{\rm GW},$ (19)
${\mathcal{A}}$ $=-8.64\times 10^{-16}\frac{\xi}{1.0}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.25}\right)^{-2}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$
$n_{\rm g}$ $=4.$
Plugging $\mathcal{A}$ and $n_{\rm g}$ into Eq. (B) in the circular limit gives us $\Delta\psi_{\rm mig}$ in Eq. (10).
C.2 Mass accretion
Assuming both companion BHs are accreting at the same constant ${\rm f}_{\rm Edd}$ during the observation window, we can write down the mass evolution as
$\dot{M}$ $={\rm f}_{\rm Edd}{\dot{M}}_{\rm Edd}=2.22\times 10^{-8}\frac{{\rm f}_{\rm Edd}}{\tau}\frac{0.1}{\epsilon}M_{z},$
$\implies M_{z}$ $=M_{z,0}\exp{\left(2.22\times 10^{-8}{\rm f}_{\rm Edd}\frac{0.1}{\epsilon}t\right)},$ (20)
where $M_{0}$ is the initial mass.
Under the SPA, a phase shift due to mass accretion in the circular limit can be expressed as (Caputo et al., 2020):
$\Delta\psi_{\rm acc}=\psi^{(0)}_{\rm TF2}\frac{25}{256}\frac{{\rm f}_{\rm Edd}}{\tau}\frac{0.1}{\epsilon}\frac{GM_{z}\eta^{-1}}{c^{3}}\left(\frac{1}{3}v_{0}^{-8}-\frac{15}{26}v^{-8}\right),$ (21)
where we have absorbed terms independent of $f$ into $\phi_{c}$ and terms proportional to $f$ in $t_{c}$. We have replaced $M_{z,0}$ with $M_{z}$ after computations, since the LISA observation duration we consider is at maximum $4~{}{\rm years}$, implying $2.22\times 10^{-8}{\rm f}_{\rm Edd}t\ll 1$ in Eq. (C.2). $\Delta\psi_{\rm acc}$ can be re-expressed as
$\Delta\psi_{\rm acc}$ $=\psi^{(0)}_{\rm TF2}\mathcal{\hat{C}}_{\rm acc}+\psi^{(0)}_{\rm TF2}C_{\rm acc}v^{-8},$ (22)
$\mathcal{\hat{C}}_{\rm acc}$ $=4.52\times 10^{-17}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.25}\right)^{-1}\frac{M_{z}}{10^{5}{\rm M}_{\sun}}\left(\frac{a_{0}}{GM/c^{2}}\right)^{4},$
$C_{\rm acc}$ $=-7.82\times 10^{-17}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.25}\right)^{-1}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$
where $a_{0}$ is the initial separation, which cannot be more than $\sim\mathcal{O}(100GM_{z}/c^{2})$ to have the MBHB merge within the LISA observation window. Therefore, $C_{I}$ is at maximum $\mathcal{O}(10^{-6})$ even for optimistic accretion rates (${\rm f}_{\rm Edd}$ $\lesssim 100$) for near equal-mass MBHBs. Hence, the first term in Eq. (22) can be dropped due to being negligible with respect to the leading-order SPA phase contribution $\psi^{(0)}_{\rm TF2}$ and we get the accretion dephasing in Eq. (22).
Appendix D Posteriors for free environmental power-law
In Fig. 8, we show posteriors for our fiducial system when the environmental power-law exponent, $n_{\rm g}$, is also a free parameter.
Figure 8: Same as in Fig. 5 but for free $n_{\rm g}$.
@@ Appendix E Bias due to ignoring gas-perturbation or eccentricity
This way of computation should also minimize statistical uncertainties in the two models during PTMCMC. Moreover, we take average of biases from two sets of independent runs to reduce statistical variance even further. We show results in the following tables for the same systems of interest that were used in Tables 3,  4,  5, and  6, respectively. The stronger that either gas-amplitude or initial eccentricity are, the larger the bias induced in the intrinsic-merger parameters.
$\boldsymbol{C_{g,\rm inj}}$ $\boldsymbol{\Delta M_{z}}$ $\boldsymbol{\Delta q}$ $\boldsymbol{\Delta\chi_{1}}$ $\boldsymbol{\Delta\chi_{2}}$ $\boldsymbol{\Delta t_{c}}$
$-10^{3}$ $-6.1$ $-6.2$ $-0.3$ $2.0$ $-4.2$
$-10^{2}$ $-1.2$ $-1.2$ $-0.2$ $0.5$ $-0.9$
$-10^{1}$ $-0.1$ $-0.1$ $0.0$ $0.0$ $-0.0$
$10^{1}$ $-0.9$ $-0.9$ $-0.3$ $0.5$ $0.7$
$10^{2}$ $-5.0$ $-5.0$ $-1.9$ $3.0$ $-4.3$
$10^{3}$ $-11.3$ $-11.4$ $-2.3$ $5.3$ $-8.9$
@@
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{\Delta M_{z}}$ $\boldsymbol{\Delta q}$ $\boldsymbol{\Delta\chi_{1}}$ $\boldsymbol{\Delta\chi_{2}}$ $\boldsymbol{\Delta t_{c}}$
$-2.5$ $0.4$ $0.4$ $0.2$ $-0.3$ $0.4$
$-2$ $2.1$ $2.1$ $0.2$ $-0.8$ $1.5$
$-1.5$ $9.1$ $9.2$ $0.2$ $-2.9$ $6.2$
@@
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}}$ $\boldsymbol{\Delta M_{z}}$ $\boldsymbol{\Delta q}$ $\boldsymbol{\Delta\chi_{1}}$ $\boldsymbol{\Delta\chi_{2}}$ $\boldsymbol{\Delta t_{c}}$
$-2$ $-10^{3}$ $5.8$ $5.8$ $0.4$ $-2.0$ $-3.9$
$-2$ $-10^{2}$ $0.9$ $0.9$ $-0.1$ $-0.2$ $0.6$
$-2$ $10^{2}$ $-1.1$ $-1.1$ $-0.0$ $0.4$ $-0.8$
$-2$ $10^{3}$ $6.5$ $6.5$ $0.8$ $-2.6$ $4.6$
@@
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}}$ $\boldsymbol{\Delta M_{z}}$ $\boldsymbol{\Delta q}$ $\boldsymbol{\Delta\chi_{1}}$ $\boldsymbol{\Delta\chi_{2}}$ $\boldsymbol{\Delta t_{c}}$
$-2$ $-10^{3}$ $1.8$ $1.8$ $0.4$ $-0.9$ $1.5$
$-2$ $-10^{2}$ $1.7$ $1.7$ $0.1$ $-0.7$ $1.2$
$-2$ $10^{2}$ $1.3$ $1.4$ $0.1$ $-0.6$ $1.0$
$-2$ $10^{3}$ $1.9$ $1.9$ $0.5$ $-1.1$ $1.6$

### v2

@@ 9 Conclusion
In this paper, we considered GWs from eccentric MBHBs embedded in a CBD to estimate if both eccentricity and gas imprints (migration and accretion) could be concurrently measured from the emitted gravitational waves observed by LISA. We study systems of interest at redshift $z=1$ with highly spinning ($\chi_{1,2}=0.9$) individual BH masses, $M_{z}$, between $10^{4}$$10^{6}~{}{\rm M}_{\sun}$ and a primary-to-secondary mass ratio, $q\in[1.2,8]$, such that the MBHBs spend at least four years in the LISA band before merging. We considered both one-year and four-year times of coalescence ($t_{c}$) to study the measurability of the parameters for systems with initial eccentricity, $e_{0}$, from $10^{-3}$$10^{-1}$ and effective gas amplitude, $C_{\rm g}$, between $-10^{-12}$ and $10^{-12}$. We assumed that the power-law ($n_{\rm g}$) scaling of the environmental effect with semi-major axis is $n_{\rm g}=4$, as expected from both migration and accretion (see Sec. 3). Due to the high expected SNRs, 150similar-toabsent150$\sim$ 150∼ 1502500250025002500, we found that LISA observations should be able to place constraints on the intrinsic-inspiral variables, $e_{0}$ and $C_{\rm g}$, as well as on the intrinsic-merger binary parameters, $\{M_{z},q,\chi_{1,2},t_{c}\}$. To account for LISA’s motion around the Sun and model the time delay interferometry response, we used the lisabeta software, and included the dephasing due to gas described by Eq. (17) into the TaylorF2Ecc waveform model. We surveyed the parameter space analytically using the Fisher formalism and then studied a few cases using Bayesian inference. Finally, we assessed whether a weak environmental imprint could be confused with a small eccentric signal using GW data alone. We itemize our main findings below.
@@
  • The cross terms between eccentricity and gas are negligible (see Appendix B). Therefore, when constraining $C_{\rm g}$ and $e_{0}$ simultaneously, the relative uncertainties on $e_{0}$ and $C_{\rm g}$ are independent of their exact magnitude (see Fig. 2).
@@
  • $C_{\rm g}\gtrsim 10^{-14}$ is constrained to $<50$ per cent relative error for circular binaries and $C_{\rm g}\gtrsim 10^{-13}$ for eccentric systems. This translates to confidently measuring ${\rm f}_{\rm Edd}$ to around $0.1$ and $1$ respectively in a four-year observation window for a $M_{z}=10^{5}~{}{\rm M}_{\sun}$ and $q=8.0$ MBHB embedded in an extremely thin disc and stronger gas torque near-merger (see Fig. 3 and the connection between $C_{\rm g}$ and accretion properties in Eq. 18).
@@
  • Bayesian inference verifies the results of the Fisher formalism, i.e. posteriors overlapped with the prediction from the Fisher matrix (see Fig. 5) and peaked at the same value of $C_{\rm g}$ with or without including eccentricity (see Fig. 9).
@@
  • Keeping the environmental power-law exponent $n_{\rm g}$ free leads to non-Gaussian and biased posteriors for $\{e_{0},C_{\rm g},n_{\rm g}\}$ due to degeneracy between $C_{\rm g}$ and $n_{\rm g}$, and low SNR in the early inspiral (see Fig. 6 and Appendix C).
  • An eccentric vacuum template can mimic a circular environmental signal for a weak injected gas amplitude, $|C_{g,\rm inj}|\lesssim 10^{-14}$. However, vice versa does not hold for small $e_{0,\rm inj}\gtrsim 10^{-2.5}$ (see Tables 3 and 4).
  • An injected eccentric environmental signal could be confused with a vacuum eccentric signal for $|C_{g,\rm inj}|\lesssim 10^{-13}$ and $e_{0,\rm inj}=0.01$. Similarly, the same signal could be mimicked by a circular gas-perturbed template even for $|C_{g,\rm inj}|\sim 10^{-12}$ (see Tables 5 and 6).
@@ Acknowledgements
MG, AD, and LM acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092. ST is supported by the SNSF Ambizione Grant Number: PZ00P2-202204. We thank the anonymous referee for helpful comments that improved this work. We acknowledge John G. Baker and Sylvain Marsat for providing us access to their lisabeta software. The authors also acknowledge use of the Mathematica software (Wolfram Research Inc., 2021), NumPy (Harris et al., 2020), and inspiration drawn from the GWFAST package (Iacovelli et al., 2022) regarding the python implementation of TaylorF2Ecc.
@@ Appendix A Definition of different terms
$\eta$ Symmetric mass ratio $q/(1+q)^{2}$
$v$ Characteristic velocity $(GM_{z}\pi f/c^{3})^{\frac{1}{3}}$
$\psi^{(0)}_{\rm TF2}$ leading-order circular phase contribution $(3/128\eta)v^{-5}$
$\Delta\psi_{\rm gas}$ leading-order gas-induced phasing correction in Eq. (17)
Intrinsic-merger parameters $\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$
Intrinsic-inspiral parameters $\{e_{0},C_{\rm g},n_{\rm g}\}$
Extrinsic parameters $\{D_{\rm L},\imath,\phi_{c},\lambda,\beta,\psi\}$
Fiducial parameters in the LISA frame $M_{z}=10^{5}~{}{\rm M}_{\sun},q=8,\chi_{1,2}=0.9,t_{c}=4~{}{\rm years},e_{0}=0.1,C_{\rm g}=10^{-13},n_{\rm g}=4,D_{\rm L}=6791.3~{}{\rm Mpc},\{\imath,\phi_{c},\lambda,\beta,\psi\}=0.5~{}{\rm radians}$
@@ Appendix B Gas-eccentricity cross terms?
Both $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are suppressed by powers of $e_{0}(v_{0}/v)^{19/6}$ and multiplied by ${\mathcal{A}}$, which is itself extremely small ($\sim\mathcal{O}(10^{-15})$) as per Eq. (12). Therefore, measurements of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ either requires them to be extremely large, or we need to observe the signal at a much lower frequency than what LISA can measure. Current simulations suggest that $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are at most $\mathcal{O}(10^{2})$ as per Table 8.161616In this work we focus only on the orbital-averaged value of the torque and neglect fluctuations that can have a higher magnitude (Zwick et al., 2022). Hence, we drop the cross-terms in Eq. (13) with $e_{0}(v_{0}/v)^{19/6}$ to recover the circular limit in Eq. (13).
Appendix C $C_{\rm g}$-$n_{\rm g}$ degeneracy
In Fig. 6, we saw that the marginal posteriors on both environmental amplitude and power-law did not peak at their injected values, despite using a zero noise realization. This is due to a degeneracy between the two environmental parameters, $C_{\rm g}$ and $n_{\rm g}$, coupled with the fact that the SNR at $z=1$ is insufficient to break it. To illustrate these points more clearly, we compute mismatches between two waveforms in Fig. 8: an injected waveform ($h_{\rm inj}$) with $e_{0,~{}\rm inj}=0.1$, $C_{\rm g,inj}=10^{-13}$, and $n_{\rm g,inj}=4$, and a template waveform ($h_{\rm tmp}$) with $C_{\rm g,tmp}=1.5\times 10^{-12}$ and varying $e_{0,\rm tmp}$ and $n_{\rm g,tmp}$, with the rest of the parameters set to their fiducial values for both waveforms. $C_{\rm g,tmp}$ is the same as the recovered value in Fig. 6.
Figure 8: Mismatches between injected and template waveforms as a function of different template eccentricities ($e_{0,\rm tmp}$) and environmental power-laws ($n_{\rm g,tmp}$) with injected intrinsic-inspiral parameters $e_{0,~{}\rm inj}=0.1$, $C_{\rm g,inj}=10^{-13}$, and $n_{\rm g,inj}=4$ and a template environmental amplitude $C_{\rm g,tmp}=1.5\times 10^{-12}$. We vary $e_{0,\rm tmp}$ in the range $[0.0995,0.1005]$ and $n_{\rm g,tmp}$ between $[3,5]$, and mark the injected values with a red cross marker.
The minimum SNR for which LISA could distinguish between these waveforms with more than $90$ per cent confidence can be computed based on the following criterion for eight free intrinsic parameters (Baird et al., 2013):
${\rm SNR}_{\rm min}^{2}=\frac{6.68}{{\rm Mismatch}(h_{\rm inj},h_{\rm tmp})},$ (25)
Given that the event SNR for our fiducial parameters is $\sim 378$, Eq. (25) implies that a mismatch of $\geq 10^{-4.33}$ is required to distinguish between the injected and template waveforms at $z=1$. However, we find a few combinations of $e_{0,\rm tmp}$ and $n_{\rm g,tmp}$, where the mismatch is lower than that, especially for $e_{0,\rm tmp}>0.1$ and $n_{\rm g,tmp}<3.5$ in Fig. 8. This implies that the SNR at $z=1$ is not enough to distinguish between the two waveforms for those combinations. As the prior volume occupied by the secondary modes is comparatively larger, the PTMCMC sampler correctly preferred those template values over the injected parameters, which led to the wrong recovered values. This behavior is not there if we have a higher SNR, as shown in Fig. 10.
Appendix D Some interesting posteriors
We compare posteriors on the environmental amplitude for a circular system to those for two eccentric systems in Fig. 9. The $C_{\rm g}$ posteriors in all cases peak around the injected value of $10^{-13}$. Also, as expected from Fig. 2, the shape of the posteriors for both non-zero eccentricities are almost the same. In the circular case, we have no support for $C_{\rm g}=0$. The broadening of the posterior is dominated by the extra degree of freedom in the model, rather than by the the presence of eccentricity in the signal itself.
In Fig. 10, we show posteriors for $e_{0}$, $C_{\rm g}$, and $n_{\rm g}$ at $z=0.01$ to show that high SNR leads to Gaussian posteriors in comparison to non-Gaussian ones in Fig. 6,
Figure 10: Same as in Fig. 6 but at $z=0.01$ instead of $z=1$.
@@ Appendix E Bias due to ignoring gas-perturbation or eccentricity
This way of computation should also minimize statistical uncertainties in the two models during PTMCMC. Moreover, we take average of biases from two sets of independent runs to reduce statistical variance even further. We show results in the following tables 9,  10,  11, and  12 for the same systems of interest that were used in Tables 3,  4,  5, and  6, respectively. The stronger that either gas-amplitude or initial eccentricity are, the larger the bias induced in the intrinsic-merger parameters.
$\boldsymbol{C_{g,\rm inj}[10^{-15}]}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta M_{z}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta q}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{1}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{2}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta t_{c}}}$
$-10^{3}$ $-5.8$ $-5.9$ $0.1$ $1.7$ $-3.9$
$-10^{2}$ $-1.7$ $-1.8$ $-0.3$ $0.8$ $-1.3$
$-10^{1}$ $-0.1$ $-0.1$ $0.1$ $-0.0$ $0.0$
$10^{1}$ $-1.5$ $-1.5$ $-0.5$ $0.8$ $-1.2$
$10^{2}$ $-5.1$ $-5.2$ $-1.5$ $2.7$ $-4.2$
$10^{3}$ $-22.6$ $-22.8$ $-9.2$ $14.7$ $-20.7$
@@
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta M_{z}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta q}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{1}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{2}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta t_{c}}}$
$-2.5$ $0.0$ $0.0$ $0.0$ $-0.0$ $0.0$
$-2$ $2.1$ $2.1$ $0.3$ $-0.9$ $1.6$
$-1.5$ $9.2$ $9.3$ $0.3$ $-3.1$ $6.4$
@@
$\boldsymbol{e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}[10^{-15}]}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta M_{z}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta q}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{1}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{2}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta t_{c}}}$
$10^{-2}$ $-10^{3}$ $-4.5$ $-4.5$ $0.2$ $1.5$ $-3.2$
$10^{-2}$ $-10^{2}$ $-1.2$ $-1.2$ $-0.3$ $0.7$ $-1.0$
$10^{-2}$ $10^{2}$ $1.4$ $1.4$ $0.3$ $-0.8$ $1.2$
$10^{-2}$ $10^{3}$ $-9.7$ $-9.8$ $-2.0$ $5.2$ $-8.2$
@@
$\boldsymbol{e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}[10^{-15}]}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta M_{z}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta q}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{1}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta\chi_{2}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Delta t_{c}}}$
$10^{-2}$ $-10^{3}$ $1.8$ $1.9$ $0.5$ $-1.0$ $1.6$
$10^{-2}$ $-10^{2}$ $1.9$ $1.9$ $0.2$ $-0.9$ $1.5$
$10^{-2}$ $10^{2}$ $1.7$ $1.7$ $0.3$ $-0.8$ $1.3$
$10^{-2}$ $10^{3}$ $1.7$ $1.7$ $0.1$ $-0.7$ $1.3$